# Deterministic Reversible Data Augmentation for Neural Machine Translation

### v1

@@ Abstract
Data augmentation is an effective way to diversify corpora in machine translation, but previous methods may introduce semantic inconsistency between original and augmented data because of irreversible operations and random subword sampling procedures. To generate both symbolically diverse and semantically consistent augmentation data, we propose Deterministic Reversible Data Augmentation (DRDA), a simple but effective data augmentation method for neural machine translation. DRDA adopts deterministic segmentations and reversible operations to generate multi-granularity subword representations and pulls them closer together with multi-view techniques. With no extra corpora or model changes required, DRDA outperforms strong baselines on several translation tasks with a clear margin (up to 4.3 BLEU gain over Transformer) and exhibits good robustness in noisy, low-resource, and cross-domain datasets. 111The relevant code is available at https://github.com/BITHLP/DRDA.
@@ 5 Experiments
We evaluate DRDA with translation tasks in different language pairs and translation directions to show its universal property regardless of language features. We also conduct experiments on extremely low resources and noisy scenarios to show the robustness of DRDA.222Further setup details about dataset split, preprocessing, models, and evaluation are listed in Appendix A.
@@ 5.1 Experimental Setup
Our experiments are conducted on different datasets, as detailed in Table 1. We experiment on a low resource setting with IWSLT datasets, including IWSLT14 EnDe, EnEs, and IWSLT17 EnZh, EnFr. We use larger WMT14 EnDe as a high-resource scenario dataset. The performance in extremely low resource scenarios is explored with the TED EnSk dataset. Following previous work Vaswani et al. (2017), we lowercase words in IWSLT EnDe, while keeping other datasets cased.333En, De, Fr, Zh, Es, Sk stand for English, German, French, Chinese, Spanish, and Slovak respectively.
@@
We evaluate the performance of NMT systems using BLEU. To compare with previous work Vaswani et al. (2017); Kambhatla et al. (2022b), we apply multi-bleu with multi_bleu.perl 444mosesdecoder/scripts/generic/multi-bleu.perl for IWSLT EnDe, WMT EnDe, and TED EnSk. For WMT EnDe dataset, we additionally apply compound splitting555tensorflow/tensor2tensor/utils/get_ende_bleu.sh. All other datasets are evaluated with SacreBLEU666SacreBLEU signature: nrefs:1|case:mixed| eff:no|tok:13a|smooth:exp|version:2.2.0.
@@ 5.4 Robustness to Perturbations
We validate the robustness of DRDA on two noisy datasets. The first one is IWSLT DeEn test set with synthetic perturbations. The perturbations are synthesized by traversing every character excluding space and punctuation in source sentences, and applying one of the operations with probability : (1) remove the character, (2) add a random character following the character, and (3) substitute the character with a random one. The second dataset is himl test set777https://www.himl.eu/test-sets, which contains health information and scientific summaries and differs considerably from the IWSLT training set. Cross-domain datasets have different subword distributions, and the difference can be viewed as a natural noise. The results of the noisy test sets are shown in Table 6.
@@ A.1 Datasets and Preprocessing
For IWSLT EnDe, following previous works, data is obtained with fairseq scripts888fairseq/example/translation//prepare_iwslt14.sh, which performs clean-corpus-n 999mosesdecoder/scripts/training/clean-corpus-n.perl with , and .
@@
We use t2t-datagen 101010tensorflow/tensor2tensor/bin/t2t-datagen script to generate WMT data, and performs clean-corpus-n with and , removing about 1% training sentence pairs. Following previous works, we validate on newstest2013 and test on newstest2014.
@@ A.4 Baseline Implementation
We use the task and loss module from the official open-source repository 111111https://github.com/dropreg/R-Drop to implement R-Drop. Weight is set to 5.
@@
CipherDAaug models are reimplemented on top of the official open-source code121212https://github.com/protonish/cipherdaug-nmt. Following their instructions, we adopt 2 keys, and set agreement loss weight .

### v2

@@ Abstract
Data augmentation is an effective way to diversify corpora in machine translation, but previous methods may introduce semantic inconsistency between original and augmented data because of irreversible operations and random subword sampling procedures. To generate both symbolically diverse and semantically consistent augmentation data, we propose Deterministic Reversible Data Augmentation (DRDA), a simple but effective data augmentation method for neural machine translation. DRDA adopts deterministic segmentations and reversible operations to generate multi-granularity subword representations and pulls them closer together with multi-view techniques. With no extra corpora or model changes required, DRDA outperforms strong baselines on several translation tasks with a clear margin (up to 4.3 BLEU gain over Transformer) and exhibits good robustness in noisy, low-resource, and cross-domain datasets. The relevant code is available at https://github.com/BITHLP/DRDA.
@@ 5 Experiments
We evaluate DRDA with translation tasks in different language pairs and translation directions to show its universal property regardless of language features. We also conduct experiments on extremely low resources and noisy scenarios to show the robustness of DRDA.111Further setup details about dataset split, preprocessing, models, and evaluation are listed in Appendix A.
@@ 5.1 Experimental Setup
Our experiments are conducted on different datasets, as detailed in Table 1. We experiment on a low resource setting with IWSLT datasets, including IWSLT14 EnDe, EnEs, and IWSLT17 EnZh, EnFr. We use larger WMT14 EnDe as a high-resource scenario dataset. The performance in extremely low resource scenarios is explored with the TED EnSk dataset. Following previous work Vaswani et al. (2017), we lowercase words in IWSLT EnDe, while keeping other datasets cased.222En, De, Fr, Zh, Es, Sk stand for English, German, French, Chinese, Spanish, and Slovak respectively.
@@
We evaluate the performance of NMT systems using BLEU. To compare with previous work Vaswani et al. (2017); Kambhatla et al. (2022b), we apply multi-bleu with multi_bleu.perl 333mosesdecoder/scripts/generic/multi-bleu.perl for IWSLT EnDe, WMT EnDe, and TED EnSk. For WMT EnDe dataset, we additionally apply compound splitting444tensorflow/tensor2tensor/utils/get_ende_bleu.sh. All other datasets are evaluated with SacreBLEU555SacreBLEU signature: nrefs:1|case:mixed| eff:no|tok:13a|smooth:exp|version:2.2.0.
@@ 5.4 Robustness to Perturbations
We validate the robustness of DRDA on two noisy datasets. The first one is IWSLT DeEn test set with synthetic perturbations. The perturbations are synthesized by traversing every character excluding space and punctuation in source sentences, and applying one of the operations with probability : (1) remove the character, (2) add a random character following the character, and (3) substitute the character with a random one. The second dataset is himl test set666https://www.himl.eu/test-sets, which contains health information and scientific summaries and differs considerably from the IWSLT training set. Cross-domain datasets have different subword distributions, and the difference can be viewed as a natural noise. The results of the noisy test sets are shown in Table 6.
@@ A.1 Datasets and Preprocessing
For IWSLT EnDe, following previous works, data is obtained with fairseq scripts777fairseq/example/translation//prepare_iwslt14.sh, which performs clean-corpus-n 888mosesdecoder/scripts/training/clean-corpus-n.perl with , and .
@@
We use t2t-datagen 999tensorflow/tensor2tensor/bin/t2t-datagen script to generate WMT data, and performs clean-corpus-n with and , removing about 1% training sentence pairs. Following previous works, we validate on newstest2013 and test on newstest2014.
@@ A.4 Baseline Implementation
We use the task and loss module from the official open-source repository 101010https://github.com/dropreg/R-Drop to implement R-Drop. Weight is set to 5.
@@
CipherDAaug models are reimplemented on top of the official open-source code111111https://github.com/protonish/cipherdaug-nmt. Following their instructions, we adopt 2 keys, and set agreement loss weight .